# Spatial-Temporal Large Language Model for Traffic Prediction

### 18 Jan 2024 v1

@@ 1 Introduction
Traffic prediction, which aims to predict future traffic features like flow and speed at specific locations using historical data, is a crucial component for intelligent transportation systems Li et al. (2023); Jiang et al. (2023); Chang et al. (2023); Li et al. (2023); Gong et al. (2023). This prediction is instrumental in optimizing traffic management Miao et al. (2024); Zhou et al. (2024) and scheduling public transportation Jin et al. (2023a); Wang et al. (2020). For instance, predicting vehicle speed aids the transportation department in optimizing traffic management. Similarly, traffic flow forecasting is vital for vehicle companies, as it enables them to efficiently allocate and schedule vehicles to satisfy expected demand.
@@ 2.2 Traffic Prediction
Traffic prediction aims to predict future traffic flow or speed based on historical traffic data, which is crucial for intelligent transportation systems Ye et al. (2021); Shao et al. (2022); Miao et al. (2023). Traditionally, the traffic prediction model is based on statistics, such as ARIMA and Kalman filter Kumar and Vanajakshi (2015); Chang et al. (2023). However, these models do not perform well due to the inherent spatial-temporal dependencies of traffic data. Later, numerous efforts have been dedicated to advancing traffic prediction techniques through the development of various neural network-based models. First, convolutional neural networks (CNNs) were applied to traffic data to capture spatial dependencies in the data Shen et al. (2018); Yuan et al. (2018). Given that CNNs are primarily designed for regular, grid-like urban areas, they encounter challenges when dealing with the non-Euclidean spatial structure of traffic data. Then, graph convolutional network (GCN) based models became popular due to their permutation-invariance, local connectivity, and compositionally Wu et al. (2019); Bai et al. (2020). More recently, attention-based models have emerged as a dominant trend Guo et al. (2019); Zheng et al. (2020); Guo et al. (2022); Jiang et al. (2023). Without taking the adjacency matrix into account, attention-based models can still model dynamic spatial correlation more effectively than GCN-based models Liu et al. (2023a). However, the structures of these models are becoming increasingly sophisticated.
@@ 3 Problem Definition
Definition 3.1 (Traffic Feature). We denote the traffic data as a tensor , where is the number of timesteps, is the number of spatial stations, and is the feature. For example, represents the traffic flow or speed.
@@ 4.3 Partially Frozen Attention (PFA) LLM
Type | Dataset | # Station (N) | # Timestep | Interval |
Flow | NYCTaxi | 266 | 4,368 | 30min |
CHBike | 250 | 4,368 | 30min | |
Speed | METR-LA | 207 | 34,272 | 5min |
PEMS-BAY | 325 | 52,116 | 5min |
@@ 5.1 Datasets
This section details the datasets employed to examine the predictive performance of the ST-LLM and baselines, with real-world traffic data from NYCtaxi Ye et al. (2021), CHBike Jiang et al. (2023), METR-LA, and PEMS-BAY Shao et al. (2022). Notably, NYCtaxi and CHBike datasets encompass two types of traffic flow data: pick-up and drop-off. Meanwhile, METR-LA and PEMS-BAY datasets focus on traffic speed. Details are as follows and the statistics of these datasets are provided in Table 1.
Method | METR-LA | PEMS-BAY | ||||
Metric | MAE | RMSE | MAPE | MAE | RMSE | MAPE |
DCRNN | 3.60 | 7.59 | 10.50% | 2.07 | 4.74 | 4.90% |
STGCN | 4.59 | 9.40 | 12.70% | 2.49 | 5.69 | 5.79% |
ASTGCN | 6.51 | 12.52 | 11.64% | 2.61 | 5.42 | 6.00% |
GWN | 3.53 | 7.37 | 10.01% | 1.95 | 4.52 | 4.63% |
AGCRN | 3.59 | 7.45 | 10.47% | 1.94 | 4.50 | 4.55% |
GMAN | 3.44 | 7.35 | 10.07% | 1.86 | 4.32 | 4.37% |
ASTGNN | 6.14 | 7.24 | 8.76% | 2.28 | 2.97 | 3.68% |
STG-NCDE | 6.35 | 14.94 | 13.49% | 2.03 | 4.58 | 4.82% |
DGCRN | 3.44 | 7.19 | 9.73% | 1.89 | 4.42 | 4.43% |
OFA | 3.17 | 6.43 | 8.55% | 1.68 | 3.71 | 3.69% |
GATGPT | 3.18 | 6.58 | 8.52% | 1.66 | 3.75 | 3.67% |
GCNGPT | 3.26 | 6.51 | 8.81% | 1.67 | 3.83 | 3.78% |
LLAMA2 | 3.37 | 6.50 | 9.12% | 1.72 | 3.69 | 3.84% |
ST-LLM | 3.09 | 6.30 | 8.41% | 1.61 | 3.61 | 3.59% |
@@ 5.3 Implementations
Aligning with contemporary practices, we divided the NYCTaxi and CHBike datasets into training, validation, and test sets using a 6:2:2 ratio, and the PEMS-BAY and METR-LA datasets into a 7:1:2 ratio. We set the historical timesteps and the future timesteps to 12 each, enabling multi-step traffic prediction. is set at 7 to represent a week’s seven days. For the traffic flow dataset, is 48, with each timestep spanning 30 minutes, while for the traffic speed dataset, is 288, corresponding to 5-minute intervals. The experiments were carried out on a system incorporating NVIDIA A100 GPUs, each with 40GB of memory. For training LLM-based models, we used the Ranger optimizer with a learning rate of 0.001, while GCN and attention-based models employed the Adam optimizer, also set at a 0.001 learning rate. The LLMs used are GPT2 and LLAMA2 7B. We configured GPT2 with six layers Zhou et al. (2023), and LLAMA2 with eight layers Jin et al. (2023b). The experiments were conducted with a batch size of 64 over 200 training epochs. Note that the experimental results are averaged across all prediction timesteps.
@@ 5.5 Results and Analysis
The comparison results with baselines are shown in Tables 2 and 3. We can make the following observations. (1) LLM-based methods yield superior prediction results, especially in traffic speed prediction scenarios with larger data volumes, with ST-LLM exhibiting the most effective performance. The ST-LLM outperforms other LLMs in six traffic prediction scenarios, demonstrating its superior accuracy in handling diverse traffic data across various datasets. (2) OFA and LLAMA2 are competent but surpassed by ST-LLM which achieves a 22.5% average MAE improvement over OFA and 20.8% over LLAMA2. This may be due to OFA’s ineffective traffic data embedding, making it difficult for LLMs to understand the spatial-temporal dependencies between data. Despite LLAMA2’s larger size and complexity, it doesn’t directly translate to better traffic prediction than ST-LLM. GATGPT and GCNGPT do not extract temporal representations of traffic data to influence the LLM to capture spatial-temporal dependencies. (3) Attention-based models like ASTGNN and GMAN exhibit varied performance across different datasets. They performed quite well in some cases but were always inferior to ST-LLM. This variability could be attributed to the limitations of traditional attention mechanisms in handling complex spatial-temporal embeddings, especially when compared to large language models. (4) GNN-based Models such as GWN and DGCRN demonstrate competitive performance, particularly in specific metrics, but still cannot outperform ST-LLM. This suggests that while GNNs effectively capture spatial dependencies, their temporal analysis capabilities might not be as advanced as the ST-LLM, which limits their overall performance.
@@ 5.7 Parameter Analysis
Within the ST-LLM framework shown in Figure 4, represents the hyperparameter that determines the number of multi-head attention layers to unfreeze during the training process. As depicted in Figure 4 (a), an analysis of the NYCTaxi Pick-up dataset reveals that with an increase in , the WAPE initially decreases. The trend suggests that unfreezing additional layers up to a certain threshold can enhance the performance of the ST-LLM. However, this positive effect inverts when exceeds 2, at which point the model’s performance starts to degrade, hinting at the diminishing benefits associated with unfreezing more layers. In a nutshell, the optimal of taxi flow prediction is set to 2. Traffic speed prediction also has the same parameter settings. Figure 4 (b) illustrates a similar pattern on the CHBike Pick-up dataset. The MAE indicates an increase in performance as is set to 1, with the lowest MAE observed at this value. This reinforces the observation that a single unfrozen multi-head attention layer is optimal for minimizing absolute errors, and the model achieves the balance between complexity and performance. This optimal point suggests that unfreezing more layers does not contribute to improved accuracy and might even degrade the performance of ST-LLM.

### 23 Jan 2024 v2

@@ 1 Introduction
Traffic prediction, which aims to predict future traffic features like flow at specific locations using historical data, is a crucial component for intelligent transportation systems Li et al. (2023); Jiang et al. (2023); Chang et al. (2023); Li et al. (2023); Gong et al. (2023). This prediction is instrumental in optimizing traffic management Miao et al. (2024); Zhou et al. (2024) and scheduling public transportation Jin et al. (2023a); Wang et al. (2020). For instance, accurately predicting bike flow benefits the transportation department in optimizing bike management. Similarly, taxi flow forecasting is vital for vehicle companies, as it enables them to efficiently allocate and schedule vehicles to satisfy expected demand.
@@ 2.2 Traffic Prediction
Traffic prediction aims to predict future traffic features based on historical traffic data, which is crucial for intelligent transportation systems Ye et al. (2021); Shao et al. (2022); Miao et al. (2023). Traditionally, the traffic prediction model is based on statistics, such as ARIMA and Kalman filter Kumar and Vanajakshi (2015); Chang et al. (2023). However, these models do not perform well due to the inherent spatial-temporal dependencies of traffic data. Later, numerous efforts have been dedicated to advancing traffic prediction techniques through the development of various neural network-based models. First, convolutional neural networks (CNNs) were applied to traffic data to capture spatial dependencies in the data Shen et al. (2018); Yuan et al. (2018). Given that CNNs are primarily designed for regular, grid-like urban areas, they encounter challenges when dealing with the non-Euclidean spatial structure of traffic data. Then, graph convolutional network (GCN) based models became popular due to their permutation-invariance, local connectivity, and compositionally Wu et al. (2019); Bai et al. (2020). More recently, attention-based models have emerged as a dominant trend Guo et al. (2019); Zheng et al. (2020); Guo et al. (2022); Jiang et al. (2023). Without taking the adjacency matrix into account, attention-based models can still model dynamic spatial correlation more effectively than GCN-based models Liu et al. (2023a). However, the structures of these models are becoming increasingly sophisticated.
@@ 3 Problem Definition
Definition 3.1 (Traffic Feature). We denote the traffic data as a tensor , where is the number of timesteps, is the number of spatial stations, and is the feature. For example, represents the traffic pick-up or drop-off flow.
@@ 5.1 Datasets
This section details the datasets employed to examine the predictive performance of the ST-LLM and baselines, with real-world traffic data from NYCtaxi Ye et al. (2021), CHBike Jiang et al. (2023).
NYCTaxi. The NYCTaxi dataset comprises over 35 million taxi trips in New York City (NYC), systematically categorized into 266 virtual stations. Spanning three months from April 1st to June 30th, 2016, it includes 4,368 timesteps, each representing a half-hour interval.
CHBike. Consisting of approximately 2.6 million Citi bike orders, the CHBike dataset reflects the usage of the bike-sharing system in the same period as the NYCTaxi dataset, from April 1st to June 30th, 2016. After filtering out stations with few orders, it focuses on the 250 most frequented stations. The dataset aligns with NYCTaxi in terms of time, covering 4,368 timesteps with each timestep representing a 30-minute interval.
@@ 5.3 Implementations
Aligning with contemporary practices, we divided the NYCTaxi and CHBike datasets into training, validation, and test sets using a 6:2:2 ratio. We set the historical timesteps and the future timesteps to 12 each, enabling multi-step traffic prediction. is set at 7 to represent a week’s seven days. is 48, with each timestep spanning 30 minutes. The experiments were carried out on a system incorporating NVIDIA A100 GPUs, each with 40GB of memory. For training LLM-based models, we used the Ranger optimizer with a learning rate of 0.001, while GCN and attention-based models employed the Adam optimizer, also set at a 0.001 learning rate. The LLMs used are GPT2 and LLAMA2 7B. We configured GPT2 with six layers Zhou et al. (2023), and LLAMA2 with eight layers Jin et al. (2023b). The experiments were conducted with a batch size of 64 over 200 training epochs. Note that the experimental results are averaged across all prediction timesteps.
@@ 5.5 Results and Analysis
The comparison results with baselines are shown in Tables 1. We can make the following observations. (1) LLM-based methods yield superior prediction results, with ST-LLM exhibiting the most effective performance. The ST-LLM outperforms other LLMs in six traffic prediction scenarios, demonstrating its superior accuracy in handling diverse traffic data across various datasets. (2) OFA and LLAMA2 are competent but surpassed by ST-LLM which achieves a 22.5% average MAE improvement over OFA and 20.8% over LLAMA2. This may be due to OFA’s ineffective traffic data embedding, making it difficult for LLMs to understand the spatial-temporal dependencies between data. Despite LLAMA2’s larger size and complexity, it doesn’t directly translate to better traffic prediction than ST-LLM. GATGPT and GCNGPT do not extract temporal representations of traffic data to influence the LLM to capture spatial-temporal dependencies. (3) Attention-based models like ASTGNN and GMAN exhibit varied performance across different datasets. They performed quite well in some cases but were always inferior to ST-LLM. This variability could be attributed to the limitations of traditional attention mechanisms in handling complex spatial-temporal embeddings, especially when compared to large language models. (4) GNN-based Models such as GWN and DGCRN demonstrate competitive performance, particularly in specific metrics, but still cannot outperform ST-LLM. This suggests that while GNNs effectively capture spatial dependencies, their temporal analysis capabilities might not be as advanced as the ST-LLM, which limits their overall performance.
@@ 5.7 Parameter Analysis
Within the ST-LLM framework shown in Figure 4, represents the hyperparameter that determines the number of multi-head attention layers to unfreeze during the training process. As depicted in Figure 4 (a), an analysis of the NYCTaxi Pick-up dataset reveals that with an increase in , the WAPE initially decreases. The trend suggests that unfreezing additional layers up to a certain threshold can enhance the performance of the ST-LLM. However, this positive effect inverts when exceeds 2, at which point the model’s performance starts to degrade, hinting at the diminishing benefits associated with unfreezing more layers. In a nutshell, the optimal of taxi flow prediction is set to 2. Figure 4 (b) illustrates a similar pattern on the CHBike Pick-up dataset. The MAE indicates an increase in performance as is set to 1, with the lowest MAE observed at this value. This reinforces the observation that a single unfrozen multi-head attention layer is optimal for minimizing absolute errors, and the model achieves the balance between complexity and performance. This optimal point suggests that unfreezing more layers does not contribute to improved accuracy and might even degrade the performance of ST-LLM.